2.5g of P4 reacts with an excess amount of F2 according to the reaction shown below. determine the maximum amount of PF5 produced? (P: 30.97g/mol, F:19g/mol)
P4+10F2 -> 4PF5
1
Expert's answer
2020-04-13T02:25:38-0400

Solution:

P4+10F2 -> 4PF5

nPF5=4nP4 - according to stoichiometric ratios and with an excess of fluorine

nP4=mP4/MP4=2.5g/(4*30.97g*mol-1)=0.02mol

nPF5=4*0.02mol=0.08mol

mPF5=nPF5*MPF5=0.08mol*(30.97g*mol-1+5*19g*mol-1)=10.1g

Answer:

Maximum amount of PF5 - 10,1g